### v1

# Model Internal Sleuthing: Finding Lexical Identity and Inflectional Morphology in Modern Language Models

Abstract
Large transformer-based language models dominate modern NLP, yet our understanding of how they encode linguistic information is rooted in studies of early models like BERT and GPT-2. To better understand today’s language models, we investigate how both classical architectures (BERT, DeBERTa, GPT-2) and contemporary large language models (Pythia, OLMo-2, Gemma-2, Qwen2.5, Llama-3.1) represent lexical identity and inflectional morphology. We train linear and nonlinear classifiers on layer-wise activations to predict word lemmas and inflectional features. We discover that models concentrate lexical information linearly in early layers and increasingly nonlinearly in later layers, while keeping inflectional information uniformly accessible and linearly separable throughout the layers. Further analysis reveals that these models encode inflectional morphology through generalizable abstractions, but rely predominantly on memorization to encode lexical identity. Remarkably, these patterns emerge across all 16 models we test, despite differences in architecture, size, and training regime (including pretrained and instruction-tuned variants). This consistency suggests that, despite substantial advances in LLM technologies, transformer models organize linguistic information in similar ways, indicating that these properties could be fundamental for next token prediction and are learned early during pretraining. Our code is available at https://github.com/ml5885/model_internal_sleuthing
Model Internal Sleuthing: Finding Lexical Identity and Inflectional Morphology in Modern Language Models
[…]
1 Introduction
Large transformer-based language models (LMs) are widely used for tasks such as text generation, question answering, and code completion (Workshop, 2023; Groeneveld et al., 2024; Grattafiori et al., 2024; Hui et al., 2024). However, how these models internally represent linguistic information remains an active research area. Prior work suggests a hierarchical organization where different layers specialize in capturing distinct levels of linguistic structure, from surface features to syntax and semantics (Jawahar et al., 2019; Tenney et al., 2019; Rogers et al., 2020).
But these studies focus only on first-generation LMs such as BERT and GPT-2 (Devlin et al., 2019; Radford et al., 2019). Since then, language technology has transformed dramatically - today’s models are far larger, trained on much more data, and adapted through extensive post-training procedures (Brown et al., 2020; Groeneveld et al., 2024; Lambert et al., 2025). We ask: how do modern LMs encode linguistic structure, and how do these representations differ from those in earlier models?
Consider the words walk, walked, jump, and jumped. Do language models group words by shared meaning (walk, walked) or by shared grammar (walked, jumped)? More broadly, where and how do LMs encode a word’s lemma (its lexical identity) and its grammatical form (inflectional morphology)?
To answer this, we train classifiers to predict either a word’s lemma or its inflection (e.g., tense, number, degree) from hidden state activations, using annotated sentences from the Universal Dependencies GUM corpus (Zeldes, 2017). We evaluate classifiers on 16 language models spanning diverse architectures and sizes. We find that:
  1. 1.
    Lexical information is predominantly encoded in early layers, while inflectional morphology is distributed more uniformly across layers.
  2. 2.
    These layer-wise distributions hold across all models, despite large-scale changes in architecture, amount of pretraining, and post-training procedures.
  3. 3.
    The effective dimensionality of representations generally declines with layer depth, though in some models the lowest point is in the middle layers - where just one or two dimensions capture nearly all the variance.
2 Model Internal Sleuthing
Figure 1: Overview of our classifier methodology. We extract hidden state activations from each model layer for target words and train two classifiers: one predicting inflection and another predicting lemma. We evaluate three classifier architectures (linear regression, MLP, and random forest) for each task to deduce information accessibility and linear separability across the layers of the language model. Performance differences between linear and non-linear classifiers reveal the encoding structure of linguistic information.
To study how language models encode linguistic information, we train simple probabilistic classifiers on the models’ internal representations or activations. [^1] We use representations, activations, and neurons interchangeably. By observing whether classifiers can recover word-level properties like lemma or inflection from a given representation, we can infer which layers, if any, linguistic information is present, and how directly (e.g., linearly or not). High classifier performance at a given layer suggests that the relevant information is encoded and accessible at that layer. Additionally, comparing performance across layers tells us how lexical and inflectional information is distributed throughout the model.
Figure 1 illustrates our approach. For each layer of each model, we extract hidden state activations for a target word and use them as input features to train specialized classifiers that predict either inflection or lemma. In our experiments, we train three different classifiers: linear regression, two-layer multi-layer perceptron (MLP), and random forest.
2.1 Linear Regression Classifier
Consistent with best practices for probing Hewitt and Liang (2019); Liu et al. (2019), we use simple linear regression classifiers. These models are solved via closed-form ridge regression Hastie et al. (2009), using the equation:
$W=(X_{\text{train}}^{T}X_{\text{train}}+\lambda I)^{-1}X_{\text{train}}^{T}Y_{\text{train}}$ (1)
where $X_{\text{train}}\in\mathbb{R}^{m\times d}$ contains $d$-dimensional representations for each of the $m$ examples in the training set, $Y_{\text{train}}\in\mathbb{R}^{m\times c}$ is the one-hot encoded matrix of training labels for $c$ classes, $\lambda$ controls the strength of L2 regularization, and $W\in\mathbb{R}^{d\times c}$ is the learned weight matrix of the classifier. Test predictions are given by:
$\hat{Y}_{\text{test}}=X_{\text{test}}W$ (2)
2.2 Random Forest Classifier
Random forests provide a complementary approach to linear models by handling complex feature interactions and decision boundaries. Unlike linear models that assume information is linearly separable, random forests combine an often large number of decision trees via model averaging to robustly identify patterns (Breiman, 2001).
Since our classification tasks involve a large number of classes (thousands of unique lemmas and multiple inflection categories), we use a one-vs-all approach. For each class $j$, we train a binary random forest classifier that distinguishes class $j$ from all other classes. Each binary classifier outputs a probability score $s_{j}(x)\in[0,1]$ representing the confidence that input $x$ belongs to class $j$. The final prediction is assigned based on the maximum score across all classes:
$\hat{y}=\operatorname{argmax}_{j\in\{1,2,...,c\}}s_{j}(x)$ (3)
where $s_{j}(x)$ is the probability that the trained binary classifier assigns to class $j$ for the input $x$ [^2] We use scikit-learn’s implementation (Pedregosa et al., 2011) with the predict_proba() method to obtain probability scores.
2.3 MLP Classifier
To test for more general non-linearity, we train a simple two-layer MLP with ReLU activation, defined as:
$\hat{Y}=\text{softmax}(\text{ReLU}(X_{\text{train}}W_{1})W_{2})$ (4)
where $W_{1}\in\mathbb{R}^{d\times h}$ and $W_{2}\in\mathbb{R}^{h\times c}$ are the weight matrices and $h=64$ is the hidden layer size (Rosenblatt, 1958) [^3] Bias terms are omitted for brevity. Two-layer MLPs with ReLU activation are universal function approximators, capable of approximating any continuous function to arbitrary precision given sufficient width (Hornik et al., 1989).
Comparing performance across these three classifier types provides insights into how information is encoded: if linear classifiers perform comparably to more complex models, information is likely linearly encoded; if random forests or MLPs significantly outperform linear models, information is non-linearly encoded (Belinkov and Glass, 2019).
2.4 Training and Evaluation
We stratify all datasets into train, validation and test splits. We train classifiers using the training set, select hyperparameters using the validation set, and evaluate on the held-out test set using accuracy and macro F1. We solve equation 1 in closed-form to identify $W\in\mathbb{R}^{d\times c}$, tune random forests for number of trees & maximum depth, and learn $W_{1}$ and $W_{2}$ for the MLP via stochastic gradient descent using the AdamW optimizer on cross-entropy loss (Kingma and Ba, 2014) [^4] We use weight decay for regularization for MLP models.
Linear Separability Gap
To quantify whether a layer’s information is linearly encoded or not, we define the linear separability gap as the difference in accuracy between non-linear and linear models. A larger gap indicates the signal is buried in non-linear structure, while a small gap suggests it is directly accessible. In practice, this gap is typically between -0.3 and 0.3, though theoretically it can range from -1 to 1.
Model Parameters Pretraining Data
Encoder-only
BERT Base 110M 12.6B tokens11footnotemark: 1
BERT Large 340M 12.6B tokens11footnotemark: 1
DeBERTa V3 Large 418M 32B tokens11footnotemark: 1
Decoder-only
GPT 2 Small 124M 8B tokens11footnotemark: 1
GPT 2 Large 708M 8B tokens11footnotemark: 1
GPT 2 XL 1.5B 8B tokens11footnotemark: 1
Pythia 6.9b 6900M 300B tokens
Pythia 6.9b Tulu 6900M 300B tokens
OLMo 2 7B 7300M 4T tokens
OLMo 2 7B Instruct 7300M 4T tokens
Gemma 2 2B 2610M 2T tokens
Gemma 2 2B Instruct 2610M 2T tokens
Qwen2.5 1.5B 1540M 18T tokens
Qwen2.5 1.5B Instruct 1540M 18T tokens
Llama 3.1 8B 8000M 15T tokens
Llama 3.1 8B Instruct 8000M 15T tokens
Table 1: Overview of models used in this study. 1Converted from GB to tokens using the approximation that 1GB of data is approximately 200M tokens in English (Chang et al., 2024).
Figure 2: Linguistic accuracy of different classifiers across model layers. The top row shows lexeme prediction accuracy, while the bottom row shows inflection prediction accuracy. Columns represent different classifier types: Linear Regression (left), MLP (middle), and Random Forest (right). Each line represents a different model.
3 Experiments
Using the classifier methodology introduced in § 2, we describe the components of our experimental setup: the dataset, model suite, and procedure for extracting token-level representations.
3.1 Dataset
We use the Universal Dependencies Georgetown University Multilayer (GUM) Corpus in our experiments (Zeldes, 2017). This corpus provides rich annotations for English texts across various genres. We curate a dataset containing 54,816 data points from 8,415 unique sentences, with 8 inflection types, 7,848 unique lemmas, and 11,720 unique word forms.
Each data point corresponds to a target word within its sentence context and includes the full sentence text, the target word’s position in the sentence, its lemma (base form), and its inflection label (grammatical category).
Formally, a data point is a tuple $(S,i,\ell,c)$ where $S=\{w_{0},w_{1},...,w_{n}\}$ is a sentence, $i$ is the index of the target word $w_{i}$ within the sentence, $\ell$ is the lemma of $w_{i}$, and $c$ is the inflection category of $w_{i}$.
The data covers three main syntactic categories: nouns (49.5%), verbs (31.2%), and adjectives (19.4%). The most frequent inflection categories are singular nouns (36.2%), verb base forms (18.4%), and positive adjectives (18.1%), while comparative (0.7%) and superlative (0.5%) forms are less common.
Sentences are relatively concise with an average length of 6.5 words, making them ideal for our analysis as they reduce contextual complexity while providing sufficient context. We split the dataset into training, validation, and test sets with a 70-10-20 ratio using stratified sampling based on inflection labels to ensure balanced representation across all grammatical categories.  [^5] See Appendix B for complete details including tokenization statistics and visualizations
3.2 Models
We study a diverse set of pretrained transformer language models available via the Hugging Face Transformers library (Wolf et al., 2020). Our selection includes different model architectures (encoder-only, decoder-only), sizes and training regimes. Table 1 lists the models used in this study.
3.3 Representation Extraction
For each data point $(S,i,\ell,c)$, we tokenize the sentence $S$ using model-specific tokenizers and process it through the model to obtain hidden states from all layers.  [^6] For target words split into multiple subword tokens, we use the representation of the last subword token to represent the entire word (Devlin et al., 2019). These representations form the basis of our layer-wise analysis of how linguistic information is encoded throughout the model.
4 Results
Figure 3: Classifier selectivity across model layers. Selectivity is defined as the difference between linguistic and control task accuracy, with higher values indicating better generalization rather than memorization. The top row shows lexeme selectivity; the bottom row shows inflection selectivity. Columns represent different classifier types: Linear Regression (left), MLP (middle), and Random Forest (right). Each line represents a different model.
4.1 Lexical Identity
Linear Regression Classifiers
Lemma prediction accuracy using linear regression shows a consistent declining pattern across layer depth for all models, beginning with accuracy between 0.8-1.0 in early layers and declining toward final layers (Figure 2, top left). DeBERTa-v3-Large exhibits the sharpest decline (0.9 $\rightarrow$ 0.3), while Pythia-6.9B maintains the highest sustained performance in later layers. Performance varies significantly across parts of speech, with models achieving highest accuracy on frequent categories like nouns and verbs, while struggling with infrequent categories like pronouns and prepositions due to limited training examples (Table 10).
MLP Classifiers
Two-layer MLP classifiers show higher overall accuracies and less pronounced decline compared to linear regression (Figure 2, top middle). Pythia models show the strongest performance throughout, while GPT-2, BERT and Qwen2.5 variants show the largest improvements from linear to MLP classifiers. The same frequency-dependent performance patterns persist across parts of speech (Table 11).
Random Forest Classifiers
Random forest results for lemma prediction are not reported due to computational constraints arising from thousands of unique lemma classes.
4.2 Inflectional Morphology
Linear Regression Classifiers
Inflectional morphology prediction using linear regression exhibits consistently high performance (0.9-1.0) across all layers and models (Figure 2, bottom left). The accuracy curves show little degradation from early to late layers. Between classifiers trained on different language model representations, there is minimal variance in accuracy. While most morphological forms achieve high accuracy, comparative and superlative adjective forms are most challenging (Table 8).
MLP Classifiers
MLP classifiers show nearly identical patterns to linear regression, maintaining accuracy between 0.9-1.0 across all layers (Figure 2, bottom middle). This highlights the fact that inflectional morphology is encoded very linearly, as additional nonlinear capacity provides no meaningful benefit. The pattern of lower performance on comparative and superlative forms persists with MLP classifiers (Table 9).
Random Forest Classifiers
Random forest classifiers exhibit lower overall accuracy (0.65-0.90) and a declining trend across layers (Figure 2, bottom right), contrasting sharply with the flat, high-performance curves of linear regression and MLP classifiers.
Figure 4: Fraction of maximum components needed to achieve different explained variance thresholds across layers. Each subplot shows a different variance threshold (50%-100%), with lines colored by model. Note that 100% explained variance always requires the full component set. Models fall into two groups: in one, dimensionality declines gradually; in the other, the middle layers require only one component. See additional plots in Figure 8.
Model $d_{\text{model}}$ ID50 ID70 ID90
First Mid Final First Mid Final First Mid Final
BERT-Base 768 123 100 88 244 212 192 461 451 446
BERT-Large 1024 138 105 85 286 226 208 567 527 554
DeBERTa-v3-Large 1024 196 133 29 377 299 113 688 635 423
GPT-2-Small 768 37 1 1 152 1 1 402 1 3
GPT-2-Large 1280 24 1 95 172 1 284 583 1 726
GPT-2-XL 1600 113 1 118 340 1 356 838 1 914
Pythia-6.9B 4096 391 1 96 865 1 517 1952 1 1925
Pythia-6.9B-Tulu 4096 390 1 244 862 1 832 1949 1 2292
OLMo-2-7B 4096 404 310 41 833 896 299 1772 2279 1550
OLMo-2-7B-Instruct 4096 404 358 111 833 974 567 1772 2361 1964
Gemma-2-2B 2304 216 8 11 505 130 70 1129 794 611
Gemma-2-2B-Instruct 2304 222 22 8 520 198 57 1153 899 572
Qwen-2.5-1.5B 1536 184 1 9 399 1 50 835 1 452
Qwen-2.5-1.5B-Instruct 1536 184 1 11 394 1 70 820 1 533
Llama-3.1-8B 4096 373 240 35 789 727 187 1722 2051 1119
Llama-3.1-8B-Instruct 4096 372 215 31 788 664 181 1722 1957 1093
Table 2: Number of principal-component axes required to reach 50% (ID50), 70% (ID70) and 90% (ID90) explained variance in the first, middle and last layers of each model.
5 Analysis
Our experiments reveal clear patterns in how transformer models encode linguistic information. We organize our analysis around six key aspects: distinct encoding patterns, classifier complexity, generalization versus memorization, representation dimensionality, the role of model architectures & training dynamics, and the effects of tokenization.
Distinct Encoding Patterns
We find consistent and distinct encoding patterns for lexical and inflectional information. Lexical information is most accessible in early layers (0-2) and then declines, while inflectional information remains uniformly high across all layers (Figure 2). This separation suggests that lexical identity is captured as an early, surface-level feature, while morphological features remain explicit throughout. This extends the hierarchical view of transformer representations advanced by Tenney et al. (2019) and Jawahar et al. (2019), by showing that surface morphology is not relegated to early layers alone but is preserved as a linearly decodable signal at every depth.
Differences in Classifier Complexity
We find that different types of linguistic information require different levels of classifier complexity to extract (Figure 2), which we quantify using the linear separability gap from § 2. Inflectional features show a small gap across all models, indicating that they remain linearly decodable. Lexical features, by contrast, show an increasingly large gap in later layers, suggesting that lexical features become progressively more non-linearly encoded (Figure 7 in Appendix C).
Generalization vs. Memorization
A key question is whether classifiers detect genuine linguistic structure or simply memorize training examples. We implement control tasks following Hewitt and Liang (2019), where words are randomly but consistently mapped to arbitrary labels. High classifier selectivity - the accuracy difference between linguistic task accuracy and control task accuracy - indicates generalization over memorization.
As shown in Figure 3, inflection prediction shows strong selectivity (0.4-0.6), especially in later layers, suggesting that models develop generalizable abstractions for morphological patterns. In contrast, lemma prediction has near-zero selectivity across all models, layers, and classifier types, suggesting that lexical information is retrieved primarily through memorization. [^7] This is perhaps unsurprising given the nature of the lemma task: it involves thousands of unique classes, many of which appear infrequently in the training data.
This aligns with the notion that morphological rules are systematic and generalizable (e.g., adding -ed to form past tense), while lexical identity is item-specific. It also connects to previous work on the increasing contextualization of representations in later layers of transformer models (Ethayarajh, 2019). Subramani et al. (2022) similarly found that steering vectors become less effective in later layers and that representations reflect semantic similarity less in later layers.
Taken together, these suggest that abstract linguistic properties become increasingly entangled with contextual information, making them increasingly harder to extract with linear methods but still recoverable through non-linear transformations.
Representation Dimensionality
[…]
Models fall into two groups (Table 2). In one group (BERT, DeBERTa, Gemma, Llama 3.1), dimensionality declines gradually across layers. In the other (GPT-2, Qwen2.5, Pythia, OLMo-2), the middle layers require just one component to reach 50-99% variance, before rising again towards the final layers of the model
Importantly, these low-dimensional middle layers are where inflectional features are most linearly separable and where lexical features show the largest non-linear gap (see Figure 7). In other words, middle layers learn to compress representations into a tight subspace that cleanly encodes inflectional morphology but entangles lexemes. Later layers expand dimensionality again to support generation or deeper contextualization while preserving linearly accessible inflectional information.
Model Architectures
Despite differences in architecture (encoder vs. decoder), model size, and training objectives, all models show remarkably consistent patterns: relatively flat, high-accuracy curves for inflection prediction and declining accuracy for lemma prediction in later layers (Figure 2). This holds across causal decoder (GPT-2, Pythia, OLMo-2, Qwen, Gemma, Llama 3.1) and bidirectional encoder models (BERT, DeBERTa).
Surprisingly, instruction-tuned models like Qwen-2.5-Instruct, Llama-3-8B-Instruct, and Pythia-6.9B-Tulu display nearly identical trends to their base counterparts (Appendix CTable 7). This suggests that instruction tuning does not substantially alter how lexical or inflectional features are encoded.
Training Dynamics
We also examined how these patterns develop during pre-training by training classifiers at intermediate model checkpoints for both OLMo-2 and Pythia (Figure 9 and Figure 10). Lexeme classification reaches peak accuracy early in training before slowly declining, while inflectional accuracy remains high and stable across both checkpoints and layer depth. This result suggests that models establish the separation between lexical and inflectional information early and maintain this structure throughout training. Taken together with our model architecture findings, we find that separating lexical and inflectional information facilitates effective next-token prediction.
Effects of Tokenization
Figure 5: Effect of tokenization strategy on analogy completion rank. Each point corresponds to a model (color) and analogy (shape). The x-axis is the rank using whole-word representations. The y-axis is the rank using tokenized representations. Here, rank means the position of the expected word when all vocabulary words are sorted by similarity to the resulting embedding from vector arithmetic; lower is better. Points above the gray $y{=}x$ line mean tokenization hurts performance.
Tokenization is an essential component of language modeling. To test how tokenization influences our findings, we use analogy completion tasks (e.g., man:king::woman:?) and compare two approaches: averaging subtoken embeddings after standard tokenization and summing embeddings from whole-word tokens.
For each approach, we perform vector arithmetic on word representations (e.g., king - man + woman). We measure performance by ranking all vocabulary words by cosine similarity to the resulting representation, and observe how highly the expected word (e.g., queen) ranks, with a lower rank indicating better performance.
Whole-word representations markedly outperform averaged subtokens across all models (Figure 5), implying that linguistic regularities are primarily stored in whole-word embeddings rather than compositionally across subtokens. Despite tokenization effects, our classifier results show consistent patterns across models using different tokenizers (see Table 5), indicating robust encoding of lexical and morphological information.
6 Related Work
Probing for linguistic information
Probing studies use supervised classifiers to predict linguistic properties from model representations (Alain and Bengio, 2017; Adi et al., 2017). Methodological improvements include control tasks to ensure probe selectivity (Hewitt and Liang, 2019). Extensive work has established that transformer models learn hierarchical linguistic structures, with different layers specializing in different information types: lower layers capture surface features and morphology, middle layers encode syntax, and upper layers represent semantics and context (Jawahar et al., 2019; Tenney et al., 2019; Liu et al., 2019; Rogers et al., 2020). Most relevant to our work, Vulić et al. (2020) found that lexical information concentrates in lower layers, while Ethayarajh (2019) showed that representations become increasingly context-specific in higher layers.
Activation steering
Beyond probing, recent work has explored manipulating model behavior by intervening on internal representations. This includes steering vectors (Subramani et al., 2022), inference-time interventions (Li et al., 2023), representation editing (Meng et al., 2022), sparse autoencoders for feature discovery (Bricken et al., 2023), and causal mediation analysis (Vig et al., 2020). These methods complement our classification-based analysis by providing insights into the functional role of different representations. See Appendix A for more detailed discussion of these interpretability approaches.
7 Conclusion
In this work, we examine how 16 transformer models from BERT to Llama 3.1 encode lexical identity and inflectional morphology. To do this, we train linear and non-linear classifiers on layer-wise activations to predict word lemmas and inflectional features, while using control tasks to measure classifier selectivity. Our investigation reveals consistent patterns across all models: lexical information concentrates in early layers and becomes increasingly non-linear in deeper layers, while inflectional information remains linearly accessible throughout the network. Moreover, our selectivity analysis shows that models primarily memorize lexical relationships, but learn generalizable patterns for inflection.
Crucially, these encoding patterns hold across different architectures, model sizes, and training procedures. These findings suggest that this linguistic information is essential for both next token prediction during pretraining and complex instruction-following, demonstrating that despite rapid advances in language technologies, certain aspects of how these models process language remain remarkably consistent.
8 Limitations
Representation Extraction for Decoder Models
Our current approach for extracting word representations from decoder-only models uses the final subword token. This assumption is an intuitive and natural choice, but may not be optimal for all architectures and models. Future work could develop better extraction methods that account for subword tokenization effects and leverage attention patterns to create more accurate word-level representations.
Form and Function in English Inflection
English contains cases where different grammatical functions share the same surface form (e.g., infinitive and non-past verb forms). We do not explicitly look at these cases in our classification/probing experiments, but these ambiguities create opportunities to better examine how models separate form from function.
English-Only Analysis
Our analysis focuses solely on English, limiting generalization to languages with richer morphological systems, where the separation of lexical identity and inflectional morphology may differ in fundamental ways.
Indirect Nature of Classifiers
While our classifier methodology follows established best practices (Hewitt and Liang, 2019; Liu et al., 2019), we only detect correlations in hidden activations, not causal mechanisms. The increasing non-linearity of lexical encoding in deeper layers suggests that linear regression, MLP and random forest classifiers may not fully capture the complete picture of how information is organized.
Appendix A Additional Related Work
A.1 Advanced Probing Methodologies
Beyond standard linear probes, there are many sophisticated approaches to understand model representations. Amnesic probing (Elazar et al., 2021) removes specific information from representations to test whether it’s necessary for downstream tasks. Minimum description length probes (Voita and Titov, 2020) balance probe complexity with performance to avoid overfitting. Causal probing (Geiger et al., 2021) aims to establish causal rather than merely correlational relationships between representations and linguistic properties. Recently, Subramani et al. (2025) find that decoding from activations directly using the Logit Lens can be used to learn confidence estimators for tool-calling agents (nostalgebraist, 2020).
A.2 Model Manipulation and Steering
Steering vectors demonstrate that specific directions in activation space correspond to high-level behavioral changes (Subramani et al., 2022). Building on this, Panickssery et al. (2024) achieves behavioral control by adding activation differences between contrasting examples. Li et al. (2023) introduce inference-time intervention, a method that shifts model activations during inference across limited attention heads to control model behavior. While these methods operate in activation space, task vectors enable arithmetic operations on model capabilities by manipulating weight space (Ilharco et al., 2023).
A.3 Feature Discovery and Mechanistic Interpretability
Recent work has explored using sparse autoencoders to discover latent features (Cunningham et al., 2023; Bricken et al., 2023), providing clearer targets for steering and interpretation. Mechanistic interpretability approaches aim to reverse-engineer the algorithms learned by neural networks (Elhage et al., 2021). Representation editing directly modifies model weights to alter specific factual associations (Meng et al., 2022). These methods complement classification-based approaches by identifying the underlying structure of learned representations and their functional significance.
Appendix B Dataset Statistics
This section provides additional statistics and visualizations for the dataset used in our experiments. Only words containing alphabetic characters and apostrophes were considered. The data covers three main syntactic categories: nouns, verbs, and adjectives, with nouns being most frequent (49.5%). For our experimental setup, we split the dataset into training, validation, and test sets with a 70-10-20 ratio using stratified sampling based on inflection labels.
B.1 Category and Inflection Distribution
Table 3 shows the distribution of word categories in our dataset. Nouns constitute nearly half of all target words, followed by verbs and adjectives.
Count Percentage
Noun 27111 49.5
Verb 17093 31.2
Adjective 10612 19.4
Table 3: Distribution of word categories in the dataset
Table 4 presents the distribution of inflection categories across the dataset. Singular nouns are the most frequent form (36.2%), while adjectival forms like comparative and superlative are relatively rare (0.7% and 0.5% respectively).
Count Percentage
singular 19830 36.2%
base 10076 18.4%
positive 9926 18.1%
plural 7281 13.3%
past 5604 10.2%
3rd_pers 1413 2.6%
comparative 403 0.7%
superlative 283 0.5%
Table 4: Distribution of inflection categories in the dataset
B.2 Sentence Statistics
Sentences in the corpus are relatively concise, with an average length of 6.5 words and a median of 5 words. The minimum number of words is 1 and the maximum is 40. This relatively short average length is ideal for our analysis, as it reduces the complexity of contextual dependencies while still providing sufficient context for meaningful representation.
Figure 6 shows the distribution of sentence lengths.
Figure 6: Distribution of sentence lengths in the dataset
B.3 Tokenization Statistics
Model Tokenizer Type
BERT Base/Large WordPiece
DeBERTa V3 Large SentencePiece
GPT-2 variants BPE
Pythia variants BPE
OLMo 2 variants BPE (tiktoken)
Gemma 2 variants SentencePiece
Qwen 2.5 variants Byte-level BPE
Llama 3.1 variants BPE (tiktoken)
Table 5: Tokenization strategies used by different model families. BPE means byte-pair encoding.
An important consideration for our analysis is how different models tokenize the words in our dataset. Table 6 shows tokenization statistics across the models we analyze. Encoder-only models like BERT and DeBERTa tend to split words into more tokens than decoder-only models like GPT-2 and Qwen2, which may affect how information is encoded across layers.
Avg. tokens Med. tokens Max tokens Percent multitoken
Model per word per word per word
BERT-base 3.11 3.0 8.0 100.0
BERT-large 3.11 3.0 8.0 100.0
DeBERTa-v3-large 3.03 3.0 6.0 100.0
GPT-2-small 1.52 1.0 5.0 42.25
GPT-2-large 1.52 1.0 5.0 42.25
GPT-2-xl 1.52 1.0 5.0 42.25
Pythia-6.9B 1.48 1.0 5.0 39.1
Pythia-6.9B-Tulu 1.48 1.0 5.0 39.1
OLMo2-7B 1.43 1.0 4.0 35.9
OLMo2-7B-Instruct 1.43 1.0 4.0 35.9
Gemma2-2B 2.19 2.0 5.0 100.0
Gemma2-2B-it 2.19 2.0 5.0 100.0
Qwen2.5-1.5B 1.43 1.0 4.0 35.9
Qwen2.5-1.5B-Instruct 1.43 1.0 4.0 35.9
Llama-3.1-8B 2.43 2.0 5.0 100.0
Llama-3.1-8B-Instruct 2.43 2.0 5.0 100.0
Table 6: Tokenization statistics across different models. Open-vocabulary models (GPT-2, Pythia-1.4B, Qwen2) require on average only 1.4–1.5 tokens per word, splitting roughly 36–42% of words into multiple tokens. By contrast, models with closed vocabularies (Gemma-2B, BERT variants, DeBERTa-v3-large) exhibit much finer subword granularity—averaging 2.2–3.1 tokens per word and splitting every word into multiple tokens. Median tokens per word reflect this: 1 for open-vocabulary models, 2–3 for closed-vocabulary models. Maximum tokens per word range from 4 (Qwen2) up to 8 (BERT).
Appendix C Linear Separability Analysis
Lexeme Inflection
Linear MLP Linear MLP
Model Slope R2 Slope R2 Slope R2 Slope R2
BERT-Base $-0.368$ $0.994$ $-0.376$ $0.924$ $0.011$ $0.063$ $0.013$ $0.096$
BERT-Large $-0.414$ $0.977$ $-0.331$ $0.878$ $0.007$ $0.025$ $0.005$ $0.020$
DeBERTa-v3-Large $-0.615$ $0.992$ $-0.527$ $0.935$ $-0.005$ $0.018$ $-0.003$ $0.005$
GPT-2-Small $-0.261$ $0.948$ $-0.049$ $0.006$ $-0.016$ $0.076$ $-0.012$ $0.032$
GPT-2-Large $-0.303$ $0.986$ $-0.389$ $0.860$ $-0.028$ $0.208$ $-0.017$ $0.183$
GPT-2-XL $-0.282$ $0.983$ $-0.367$ $0.868$ $-0.036$ $0.329$ $-0.022$ $0.315$
Pythia-6.9B $-0.213$ $0.992$ $-0.127$ $0.056$ $-0.009$ $0.053$ $-0.007$ $0.028$
Pythia-6.9B-Tulu $-0.195$ $0.988$ $-0.111$ $0.043$ $-0.005$ $0.016$ $-0.002$ $0.003$
OLMo-2-7B $-0.153$ $0.566$ $-0.186$ $0.614$ $0.011$ $0.067$ $0.015$ $0.137$
OLMo-2-7B-Instruct $-0.157$ $0.574$ $-0.178$ $0.615$ $0.009$ $0.051$ $0.013$ $0.106$
Gemma-2-2B $-0.161$ $0.413$ $-0.198$ $0.714$ $0.003$ $0.009$ $0.004$ $0.013$
Gemma-2-2B-Instruct $-0.116$ $0.233$ $-0.207$ $0.505$ $0.016$ $0.176$ $0.014$ $0.138$
Qwen2.5-1.5B $-0.260$ $0.629$ $0.040$ $0.007$ $0.004$ $0.007$ $0.015$ $0.054$
Qwen2.5-1.5B-Instruct $-0.191$ $0.392$ $0.126$ $0.077$ $0.002$ $0.002$ $0.015$ $0.050$
Llama-3.1-8B $-0.109$ $0.188$ $-0.163$ $0.387$ $0.015$ $0.105$ $0.014$ $0.101$
Llama-3.1-8B-Instruct $-0.076$ $0.103$ $-0.121$ $0.258$ $0.018$ $0.153$ $0.016$ $0.138$
Table 7: Regression slopes and R2 values for regressing probe accuracy on layers. Negative slopes indicate decreasing accuracy with layer depth.
Model 3rd person Base Comparative Past Plural Positive Singular Superlative
(n=249) (n=1,833) (n=76) (n=1,003) (n=1,247) (n=1,785) (n=3,587) (n=52)
BERT-Base 0.960 0.965 0.817 0.967 0.983 0.946 0.971 0.759
BERT-Large 0.956 0.964 0.861 0.968 0.982 0.950 0.971 0.768
DeBERTa-v3-Large 0.938 0.974 0.831 0.961 0.986 0.954 0.977 0.706
GPT-2-Small 0.828 0.958 0.840 0.956 0.974 0.941 0.964 0.754
GPT-2-Large 0.812 0.958 0.826 0.951 0.975 0.936 0.967 0.792
GPT-2-XL 0.817 0.959 0.813 0.948 0.977 0.940 0.968 0.788
Pythia-6.9B 0.886 0.972 0.904 0.964 0.989 0.957 0.977 0.907
Pythia-6.9B-Tulu 0.899 0.973 0.909 0.967 0.989 0.956 0.976 0.910
OLMo-2-1124-7B 0.938 0.968 0.902 0.972 0.981 0.923 0.966 0.888
OLMo-2-1124-7B-Instruct 0.927 0.967 0.896 0.971 0.981 0.923 0.965 0.872
Gemma-2-2B 0.901 0.968 0.797 0.969 0.986 0.947 0.974 0.833
Gemma-2-2B-Instruct 0.913 0.966 0.863 0.973 0.988 0.938 0.972 0.872
Qwen2.5-1.5B 0.856 0.950 0.802 0.942 0.972 0.919 0.957 0.688
Qwen2.5-1.5B-Instruct 0.774 0.954 0.647 0.945 0.972 0.921 0.965 0.630
Table 8: Breakdown of inflection classification accuracy by morphological feature for each model, using linear regression classifiers. Inflections are grouped by their morphological features (e.g., Past, Plural, Comparative). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale.
Model 3rd person Base Comparative Past Plural Positive Singular Superlative
(n=249) (n=1,833) (n=76) (n=1,003) (n=1,247) (n=1,785) (n=3,587) (n=52)
BERT-Base 0.973 0.969 0.910 0.972 0.989 0.959 0.974 0.939
BERT-Large 0.967 0.970 0.910 0.973 0.988 0.961 0.975 0.931
DeBERTa-v3-Large 0.954 0.976 0.925 0.966 0.989 0.962 0.979 0.867
GPT-2-Small 0.921 0.963 0.928 0.952 0.972 0.930 0.963 0.870
GPT-2-Large 0.857 0.962 0.872 0.955 0.976 0.942 0.967 0.854
GPT-2-XL 0.921 0.963 0.928 0.952 0.972 0.930 0.963 0.870
Pythia-6.9B 0.932 0.972 0.921 0.961 0.982 0.949 0.971 0.886
Pythia-6.9B-Tulu 0.948 0.974 0.932 0.964 0.983 0.949 0.971 0.897
OLMo-2-1124-7B 0.957 0.968 0.926 0.966 0.989 0.949 0.973 0.905
OLMo-2-1124-7B-Instruct 0.939 0.967 0.903 0.967 0.988 0.949 0.973 0.873
Gemma-2-2B 0.913 0.967 0.863 0.968 0.990 0.950 0.976 0.907
Gemma-2-2B-Instruct 0.930 0.970 0.878 0.975 0.989 0.946 0.974 0.906
Qwen2.5-1.5B 0.882 0.948 0.822 0.943 0.974 0.927 0.957 0.736
Qwen2.5-1.5B-Instruct 0.808 0.953 0.697 0.947 0.974 0.930 0.965 0.682
Table 9: Breakdown of inflection classification accuracy by morphological feature for each model, using MLP classifiers. Inflections are grouped by their morphological features (e.g., Past, Plural, Comparative). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale.
Model Noun Verb Adjective Adverb Pronoun Preposition Conjunction Interjection Other
(n=1,739) (n=641) (n=641) (n=23) (n=1) (n=1) (n=1) (n=1) (n=9)
BERT-Base 0.636 0.737 0.609 0.805 0.292 0.000 0.585 0.000 0.902
BERT-Large 0.684 0.777 0.653 0.826 0.580 0.154 0.662 0.065 0.897
DeBERTa-v3-Large 0.592 0.737 0.585 0.723 0.440 0.077 0.438 0.081 0.866
GPT-2-Small 0.631 0.789 0.612 0.813 0.542 0.000 0.415 0.033 0.896
GPT-2-Large 0.691 0.810 0.688 0.847 0.853 0.174 0.267 0.115 0.912
GPT-2-XL 0.713 0.827 0.708 0.847 0.724 0.222 0.311 0.241 0.899
Pythia-6.9B 0.856 0.926 0.836 0.926 0.938 0.443 0.566 0.488 0.934
Pythia-6.9B-Tulu 0.864 0.930 0.843 0.930 0.923 0.514 0.651 0.476 0.936
OLMo-2-1124-7B 0.798 0.875 0.794 0.913 0.697 0.339 0.363 0.495 0.913
OLMo-2-1124-7B-Instruct 0.798 0.868 0.792 0.902 0.606 0.339 0.331 0.495 0.910
Gemma-2-2B 0.757 0.869 0.736 0.876 0.667 0.179 0.205 0.288 0.891
Gemma-2-2B-Instruct 0.749 0.844 0.742 0.872 0.620 0.137 0.152 0.247 0.912
Qwen2.5-1.5B 0.652 0.801 0.650 0.828 0.526 0.082 0.223 0.068 0.867
Qwen2.5-1.5B-Instruct 0.642 0.800 0.632 0.831 0.544 0.082 0.245 0.068 0.877
Llama-3-8B 0.776 0.882 0.771 0.887 0.831 0.286 0.396 0.321 0.911
Llama-3-8B-Instruct 0.796 0.892 0.788 0.896 0.908 0.300 0.443 0.357 0.917
Table 10: Breakdown of lexeme classification accuracy by Part of Speech (POS) for each model, using linear regression classifiers. Lexemes are grouped by their POS tags (e.g., Noun, Verb, Adjective). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale.
Model Noun Verb Adjective Adverb Pronoun Preposition Conjunction Interjection Other
(n=1,739) (n=641) (n=641) (n=23) (n=1) (n=1) (n=1) (n=1) (n=9)
BERT-Base 0.775 0.831 0.748 0.873 0.458 0.125 0.756 0.267 0.898
BERT-Large 0.813 0.863 0.785 0.884 0.540 0.231 0.725 0.323 0.897
DeBERTa-v3-Large 0.689 0.803 0.682 0.802 0.700 0.115 0.662 0.242 0.861
GPT-2-Small 0.678 0.792 0.665 0.765 0.042 0.000 0.610 0.000 0.830
GPT-2-Large 0.754 0.837 0.755 0.827 0.347 0.188 0.596 0.385 0.871
GPT-2-XL 0.774 0.844 0.771 0.827 0.561 0.232 0.561 0.431 0.860
Pythia-6.9B 0.774 0.856 0.768 0.862 0.554 0.229 0.528 0.310 0.868
Pythia-6.9B-Tulu 0.818 0.880 0.803 0.887 0.554 0.343 0.613 0.381 0.889
OLMo-2-1124-7B 0.818 0.877 0.828 0.896 0.727 0.290 0.734 0.505 0.885
OLMo-2-1124-7B-Instruct 0.822 0.874 0.829 0.897 0.667 0.306 0.750 0.473 0.886
Gemma-2-2B 0.763 0.860 0.763 0.881 0.574 0.125 0.443 0.182 0.880
Gemma-2-2B-Instruct 0.777 0.846 0.785 0.882 0.580 0.137 0.400 0.299 0.875
Qwen2.5-1.5B 0.747 0.838 0.742 0.811 0.228 0.131 0.628 0.164 0.857
Qwen2.5-1.5B-Instruct 0.749 0.840 0.738 0.818 0.211 0.098 0.564 0.123 0.860
Llama-3-8B 0.798 0.879 0.807 0.886 0.800 0.214 0.679 0.393 0.882
Llama-3-8B-Instruct 0.824 0.893 0.826 0.895 0.831 0.257 0.689 0.429 0.887
Table 11: Breakdown of lexeme classification accuracy by Part of Speech (POS) for each model, using Multi-Layer Perceptron (MLP) classifiers. Lexemes are grouped by their POS tags (e.g., Noun, Verb, Adjective). For each group, the reported accuracy is the average of accuracies from classifiers trained at each model layer. All accuracy values are on a 0–1 scale.
Linear Separability Gap
(a) Linear separability gap for inflection prediction
(b) Linear separability gap for lemma prediction
Figure 7: Performance advantage of MLP classifiers over linear classifiers (in percentage points) across model layers. The linear separability gap measures how much a non-linear transformation improves classifier performance compared to a simple linear mapping. For inflection prediction, the gap is consistently minimal (mostly within ±$\pm 0.02$ percentage points) and sometimes negative, indicating that inflectional features are primarily encoded in a linear fashion throughout the network. By contrast, the linear separability gap for lemma prediction is relatively large (0.1–0.3 percentage points) and positive across most models
Figure 8: Intrinsic dimensionality curves for all 10 models. Each subplot shows the relationship between the percentage of maximum PCA components (x-axis) and the percentage of explained variance (y-axis) across different layers. The color gradient from purple (early layers, 0%) to yellow (late layers, 100%) indicates the relative layer depth within each model. Models like BERT, Gemma, and Llama show similar compression patterns, while GPT-2 variants, Qwen and Pythia exhibit opposite trends in their middle layers.
Figure 9: Lexeme (top row) and inflection (bottom row) prediction accuracy across normalized layer number for OLMo-2-7B checkpoints at various pretraining steps (5k-734k steps). The full model is 928k steps. Checkpoints are on a color gradient from brightest (earliest) to darkest (latest). Early checkpoints exhibit higher lexeme accuracy than later ones, while inflectional accuracy remains flat across layers and checkpoints.
Figure 10: Lexeme (top row) and inflection (bottom row) prediction accuracy across normalized layer number for Pythia-6.9B checkpoints at various pretraining steps (1-111k steps). The full model is 143k steps. Checkpoints are on a color gradient from brightest (earliest) to darkest (latest). Lexeme accuracy declines both with deeper layers and with more training, whereas inflectional accuracy stays uniformly high across all layers and checkpoints.